You arrive at a movie theater only to notice that there’s only one person selling tickets. You watch for some time and realize that on average about 80% of the time there is no one in the line or being served. People are clearly arriving randomly but on average every 30 seconds.

The average service time in seconds is:
Answer #1

Arrival rate = λ = 60/30 = 2 / min

Let Service Rate = μ / min

Probability of no customers in the system = Po = 1 - λ/μ = 80%

=> 1 - 2/μ = 0.80

=> μ = 2/(1 - 0.80) = 10 per minute

Hence, service rate = 60/10 = 6 secs per customer